OpenStudy (anonymous):

Simplify: 7x + 4y - 3x + 10z + 5y A. 4x + 9y + 10z B. 4x + 19y C. 23xyz D. 10x + 10z + 9y @sangya21

OpenStudy (anonymous):

7x + 4y - 3x + 10z + 5y = (7-3)x + (4 +5)y + 10z = 4x + 9y + 10z